# Self-Supervised Learning based on Transformed Image Reconstruction for Equivariance-Coherent Feature Representation

### v1

@@ Abstract
The equivariant behaviour of features is essential in many computer vision tasks, yet popular self-supervised learning (SSL) methods tend to constrain equivariance by design. We propose a self-supervised learning approach where the system learns transformations independently by reconstructing images that have undergone previously unseen transformations. Specifically, the model is tasked to reconstruct intermediate transformed images, e.g. translated or rotated images, without prior knowledge of these transformations. This auxiliary task encourages the model to develop equivariance-coherent features without relying on predefined transformation rules. To this end, we apply transformations to the input image, generating an image pair, and then split the extracted features into two sets per image. One set is used with a usual SSL loss encouraging invariance, the other with our loss based on the auxiliary task to reconstruct the intermediate transformed images. Our loss and the SSL loss are linearly combined with weighted terms. Evaluating on synthetic tasks with natural images, our proposed method strongly outperforms all competitors, regardless of whether they are designed to learn equivariance. Furthermore, when trained alongside augmentation-based methods as the invariance tasks, such as iBOT or DINOv2, we successfully learn a balanced combination of invariant and equivariant features. Our approach performs strong on a rich set of realistic computer vision downstream tasks, almost always improving over all baselines.
@@ 1 Introduction
Figure 1: Overview of our proposed framework. Transformation $g$ with example of rotation with angle $\theta$. The common augmentation pipeline is as DINOv2, but on $v_{2}$, a transformation is applied along with common augmentation.
We introduce a self-supervised learning (SSL) auxiliary task based on reconstruction of transformed images that improves learning of equivariance-coherent feature representations. For a schematic overview see Figure 1.
Leveraging unannotated images for image representation learning through SSL [8, 22, 9, 24, 2, 33, 45] methods is crucial for pretraining foundation models. However, most SSL methods, including e.g. the popular DINOv2 [33], are based on surrogate tasks, supporting invariance of learned features with respect to augmentations of the inputs. In addition, the most common evaluation metric for SSL in computer vision is image classification by linear probing on ImageNet-1K [12], a task with labels being invariant under image transformations used for augmentation. Consequently, SSL methods are primarily evaluated for suitability of their invariant features. For downstream tasks relying on equivariance, this may not be the best approach.
Equivariance in feature learning ensures that a model’s learned representations remain consistent under various transformations, including 2D or 3D translations, rotations, scaling, and changes in colour or illumination. ’Equivariance’ means that the information about the transformation can still be retrieved from the feature representation. In convolutional neural networks (CNNs) [35], translation equivariance is achieved using convolutional layers. This property is useful in tasks like object detection and image segmentation where transformation awareness (e.g., estimating the position and orientation of an object) is critical.
As we show, current SSL methods exhibit suboptimal performance on tasks related to equivariance, such as the synthetic task in Figure 2, primarily due to their reliance on augmentation-invariant approaches. We heal this deficit by applying our auxiliary task to a fraction $z_{\text{equi}}$ of the features and show that this allows to improve performance consistently, not only for synthetic equivariant downstream tasks, but also even for tasks associated with feature invariance.
Our contribution
  • We introduce an augmentation-based SSL loss enhancing equivariant behavior of learned features.
  • Our approach strongly outperforms competitors on all synthetic equivariance test tasks from [36].
  • Our approach performs strong on a wide range of computer vision tasks compared to IBOT [45] and DINOv2 [33], improving performance in almost all cases and clearly improving on average, where other equivariance-enhancing approaches underperform.
2 Related work
2.1 Equivariant neural networks
Since the early days of neural networks research, the exploration of symmetries in the data has played a significant role, reduced model complexity, and improved inference quality of models [16]. One might argue that without convolutional neural networks, inherently implementing approximately translational equivariance, computer vision models could not have made this progress in the field. However, built-in permutation equivariance in transformer architectures has also been the object of study [41]. Not surprisingly, equivariance in deep learning is a vivid research field which we define into sub-categories: those that study models that inherit built-in equivariance [10, 26] and models that gain this property by experience [39, 11, 37].
2.2 Self-supervised learning
State-of-the-art self-supervised learning (SSL) methods learn feature representations by automatically labelling non-labeled data and applying supervised learning techniques. The assumption is that the learned feature representation is comprehensive enough to be used later in other tasks, denoted as downstream tasks. A large variety of different SSL methods have been proposed [22, 8, 23, 9, 42, 7, 5, 29, 40, 33] and they can be classified with different metrics. For an overview, we refer to [3]. The relevant metric for our purpose is how models react to different transformations in the input space, i.e., if they maintain the information in the feature representation or if this information is lost or partly lost. Older SSL methods proposed auxiliary tasks such as Jigsaw puzzle [32] or rotation estimation [19] that foster equivariance properties as the transformation properties need to be maintained in feature space. These methods have been overtaken by matching-type methods that present two different versions of the same semantic content to the model and motivate it to map these two versions to nearby points in feature space. Semantically different images are mapped to points far apart. This is achieved either by contrastive learning approaches or by regularisation techniques like de-correlation [42, 5] or teacher-student architectures [45, 33]. Consequently, these matching-type methods learn feature representations that suppress the differences between the versions of the same semantic image. Another state-of-the-art SSL branches are mask-based approaches that remove part of the input image and reconstruct them or a transformed version of them, either in the original image space [23, 4], or in the feature space [2]. Contrastive learning has been combined with the masked approach, but only pixel-accurate translation has been applied as augmentation [25]. As these methods do not, apart from masking or cropping, rely on other transformation, they are by construction more open for equivariance. Our approach is closely related to SIE [18], a method that combines the matching approach with an explicit model of transformations applied in the input space. Our approach can be seen as an extension which does not require knowledge about the transformation parameters, i.e. it can also be applied with respect to any transformations without the need to know them explicitly as long as one can apply them to the image. In contrast to SIE, our approach reaches state-of-the-art results.
2.3 Equivariance vs. invariance
A function, in our context represented by the deep learning model, between two spaces, the input image space in the context of computer vision tasks and the space of feature presentations, is denoted as equivariant with respect to a transformation in the input space and a corresponding transformation in the output space, if the function commutes with the transformations, i.e. it does not matter if one first applies the transformation in the input space and then the function or first the mapping to the output space and then applies the corresponding transformation there. The definition includes the identity transformation in the output space such that invariance is always equivariance. However, in the computer vision literature e.g. [39, 11, 14, 17, 34, 20, 37], as we do in this paper, invariance is often opposed to equivariance to stress that all information about the transformation in the input space is still retrievable from the output space. The term equivariance is usually considered for a set of transformations, i.e. the function is said to be equivariant with respect to this set of transformations. Moreover, the set of transformations is considered a group transformation or, even stronger, a group representation of the transformation, i.e. a linear map, in the input and output space. However, not all transformations applied in computer vision can be modelled by group transformations like elastic distortions, crop-resize operations, or non-affine perspective transformations. In addition, transformations that can theoretically be formulated as group transformations might lose the corresponding group properties in practice, e.g., an image rotated around a general angle cannot be rotated back as parts of the image get lost during the forward transformation. In this paper, we do not restrict the model to be equivariant with transformations that belong to a certain further structure but argue that equivariance is no value in itself but should serve as a property to help to learn a feature representation that contains all the information necessary for all kinds of downstream tasks. We do not restrict the transformation to form a group or require them to be linear in the feature space. Instead we motivate the model to maintain the information that is necessary to maintain input output relation such that transformed images in the input space can be retrieved by the representation irrespective if the transformation forms a linear transformation, a general group transformation, or even if the definition of equivariance is only approximately fulfilled in the feature representation. We denote this as equivariant-coherence.
3 Method
For augmentation-based SSL methods, joint embedding feature matching between two augmented images is commonly used to learn image features from unlabeled data. While these methods effectively capture invariant features, they require auxiliary tasks to learn equivariant features. Our approach draws inspiration from SIE [18], which splits features into invariant and equivariant components, using the equivariant part for auxiliary tasks to learn equivariant-coherent features. A detailed comparison between our method and SIE is presented in Section 3.1.
In our proposed framework, depicted in Figure 1, we introduce transformation $g$ as an additional augmentation applied exclusively to the second view $v_{2}\in\mathbb{R}^{B\times 3\times 224\times 224}$, where $B$ denotes the batch size. Both augmented views are processed by the encoder $f$, which shares weights during pretraining. The resulting feature representations $z\in\mathbb{R}^{B\times N\times d_{\text{patch\_embed}}}$ are then divided into two distinct sets along patch embeddings, that is, the total patch embeddings $d_{\text{patch\_embed}}=d_{\text{inv}}+d_{\text{equiv}}$, where $N$ is the number of patches and $d_{\text{patch\_embed}}$ is the dimension of patch embedding:
  • Invariant features $z_{\text{inv}}^{i}\in\mathbb{R}^{B\times N\times d_{\text{inv}}}$: These features compute the self-supervised learning (SSL) loss based on the chosen method (e.g., iBOT loss).
  • Equivariant features $z_{\text{equi}}^{i}\in\mathbb{R}^{B\times N\times d_{\text{equi}}}$: These features are concatenated to reconstruct intermediate transformed images.
We employ two independent decoders, $d_{1}$ and $d_{2}$, operating on the concatenated equivariant features to reconstruct intermediate transformed images, thereby computing the reconstruction loss $L_{\text{recon}}$.
It is combined with the SSL loss $L_{\text{SSL}}$ using a weighting hyperparameter $\lambda$:
$L_{\text{total}}=L_{\text{SSL}}+\lambda L_{\text{recon}}$ (1)
Notably, the feature split ratio $|z_{\text{equi}}|/|z_{\text{inv}}|$ between invariant and equivariant features is not necessarily one. We conduct sensitivity analyses to investigate how different splitting strategies affect equivariant feature learning.
3.1 Our Method vs. SIE [18]
In SIE, the exact transformation parameters between views are required for training and modeled as a linear mapping, which is then learned to align equivariant features across views. A key challenge in SIE is obtaining prior knowledge of these transformations, which is often impractical for natural images. To address this, we artificially introduce transformations such as rotation, color jittering, Gaussian blur, and translation, ensuring that the transformation parameters are known and allowing SIE to be applied to natural images.
In contrast, our method does not require knowledge of transformation parameters to train a model with equivariant features. Instead, we learn the natural variations between two augmented views. Incorporating spatial augmentations into the standard self-supervised learning augmentation pipeline with our approach effectively facilitates the learning of equivariant features, as demonstrated in Table 2.
3.2 Transformations
Transformations $g$ Parameters Mag. range
Rotation angle $\theta$ $[-45,45]$
Color jittering brightness $B$, contrast $C$, $[-0.4,0.4]$
saturation $S$, hue $H$ $[-0.1,0.1]$
Gaussian blur radius $\sigma$ $[0.1,2]$
Translation displacement $t_{x},t_{y}$ $[-10,10]$
SE(2) angle $\theta$ $[-45,45]$
displacement $t_{x},t_{y}$ $[-10,10]$
Table 1: Considered transformations.
Additionally, we incorporate non-geometric transformations, including color jittering and Gaussian blur. The parameter magnitude ranges for these transformations are shown in Table 1.
We designed the decoder to be as simple as possible, compare yellow box in Figure 1. Specifically, we used a single linear layer to match the input dimensions of the subsequent four convolutional layers. Our goal is not to reconstruct perfect intermediate transformed images but to develop an encoder $f$ that produces better equivariant representations. Therefore, the decoder $d$ should not interfere with the learning of equivariance-coherent features during pretraining. We implemented two independent decoders for learning the intermediate reconstruction, with the number of decoders determined by the number of intermediate images. In the sensitivity analysis, we determine the optimal number of intermediate images, finding that 2 is optimal.
4 Experimental results
As our method builds on SIE [18], SIE is our first natural baseline with synthetic evaluation tasks designed to benefit from equivariant features. To evaluate the generalization of our method against state-of-the-art approaches, we go beyond using the invariant loss function $L_{\text{SSL}}$ as in SIE and explore integrating other methods. Specifically, we incorporate co-learning with iBOT [45] and DINOv2 [33], both of which are augmentation-based techniques. They are also tested on the synthetic benchmark. Finally, we evaluate on a rich set of more realistic downstream tasks. We aim to enhance the baselines’ performance on equivariance-related tasks while preserving strong results on invariance-related benchmarks, e.g. ImageNet linear probing.
4.1 Implementation Details
Architecture. We use Vision Transformers (ViTs) [15] with different configurations, specifically ViT-S/16 and ViT-L/16, as our backbone architectures for experiments. As the different baseline methods we compare to use different representation dimensions, i.e. 8192 for iBOT, 512 for SIE, and 2048 for DINOv2, we incorporate a linear head on top of the backbone to accommodate.
Afterwards, a portion $z_{\text{equi}}$ of these features $z$ is allocated for reconstruction; we do not introduce more features than the baseline methods.
Pretraining setup. Our approach uses a baseline SSL loss $L_{\text{SSL}}$ in addition to our new component $L_{\text{recon}}$. Each of the three baseline methods come with distinct training setups. The common training configuration includes ImageNet-1K as the dataset, AdamW [31] as the optimizer, and a cosine-scheduled learning rate. For the SIE-based method, where we apply only their invariant loss as $L_{SSL}$, we pretrain ViT-S/16 for 800 epochs with a batch size of 2048. The base learning rate is set to $1e{-4}$ and is linearly scaled with the batch size: $lr=1e{-4}\times\frac{\text{batch\_size}}{256}$. For the iBOT-based approach, we pretrain ViT-S/16 for 800 epochs and ViT-L/16 for 250 epochs, both with a batch size of 1024. The learning rate follows the same linear scaling strategy, with a base learning rate of $5e{-4}$. For the DINOv2-based training, we train ViT-L/16 with 100 epochs with batch size of 2048. The base learning rate is $4e{-3}$ with warmup 10 epochs.
4.2 Synthetic tasks
Figure 2: Synthetic tasks for evaluating equivariant representations. The transformation ($g$) is applied to the original image $I$, and both the transformed and original images are processed through a pretrained encoder $f$. A lightweight MLP $h$ then predicts the parameters of the applied transformation.

### v2

@@ Abstract
Self-supervised learning (SSL) methods have achieved remarkable success in learning image representations allowing invariances in them — but therefore discarding transformation information that some computer vision tasks actually require. While recent approaches attempt to address this limitation by learning equivariant features using linear operators in feature space, they impose restrictive assumptions that constrain flexibility and generalization. We introduce a weaker definition for the transformation relation between image and feature space denoted as equivariance-coherence. We propose a novel SSL auxillary task that learns equivariance-coherent representations through intermediate transformation reconstruction, which can be integrated with existing joint embedding SSL methods. Our key idea is to reconstruct images at intermediate points along transformation paths, e.g. when training on 30° rotations, we reconstruct the 10° and 20° rotation states. Reconstructing intermediate states requires the transformation information used in augmentations, rather than suppressing it, and therefore fosters features containing the augmented transformation information. Our method decomposes feature vectors into invariant and equivariant parts, training them with standard SSL losses and reconstruction losses, respectively. We demonstrate substantial improvements on synthetic equivariance benchmarks while maintaining competitive performance on downstream tasks requiring invariant representations. The approach seamlessly integrates with existing SSL methods (iBOT, DINOv2) and consistently enhances performance across diverse tasks, including segmentation, detection, depth estimation, and video dense prediction. Our framework provides a practical way for augmenting SSL methods with equivariant capabilities while preserving invariant performance.
@@ 1 Introduction
Figure 1: Overview of our framework for equivariance-coherent feature representation. As done in many joint embedding SSL methods like DINOv2, we apply two sets of image augmentations generating two different views of the original image. On the secondary view, we additionally apply a sequence of transformations $g_{i}$ (e.g., three rotations by the angles $\frac{\varphi}{3}$$\frac{2\varphi}{3}$ and $\varphi$ as shown in the figure) to enforce equivariant structure during transformed image reconstruction.The last rotation serves as the second view for the joint embedding whereas the in-between transformed images are the targets of the reconstruction task.
Self-supervised learning (SSL) methods (Chen et al. 2020a; He et al. 2020; Chen et al. 2020b; He et al. 2022b; Assran et al. 2023; Oquab et al. 2023; Zhou et al. 2022) have become crucial for pretraining foundation models by leveraging unannotated images for representation learning. However, joint embedding SSL methods, currently one of the best performing SSL methods, face a fundamental trade-off: they are designed around surrogate tasks that promote invariance, i.e. learning features that remain unchanged under input transformations. This invariance bias is reinforced due to the most popular evaluation being linear probing on ImageNet-1K (Deng et al. 2009), where classification labels are inherently invariant to transformations used in data augmentation.
While invariance is valuable for classification tasks, many computer vision applications require equivariant features that preserve transformation information rather than discarding it. Equivariance means that when an input undergoes a transformation (e.g. rotation, translation, scaling), the learned representation changes in a predictable, recoverable way. This property is essential for dense prediction tasks like object detection, segmentation, or pose estimation, where knowing precise object positions and orientations is critical for understanding where objects are in a scene. Recent methods like the Split Invariant–Equivariant (SIE) framework (Garrido et al. 2023b) attempt to address this limitation by learning both invariant and equivariant features through transformation-conditioned linear operators in latent space. While effective, SIE imposes restrictive architectural constraints and assumes equivariant mappings to be linear. Such assumptions may be too strict for complex transformations or non-rigid deformations.
Our key idea is to use augmentation information in an auxiliary task, rather than modelling equivariance in latent space.
Our method reconstructs images at intermediate points along transformation paths. For example, when training on a 30° rotation, we reconstruct images at 10° and 20° rotation angles. This intermediate reconstruction pushes the model to provide features containing information on the transformation and thus encourages equivariant learning. Compared to SIE, our approach offers two key advantages: (1) it removes linearity constraints on equivariant mappings, broadening the function space for learning unseen or complex transformations, (2) it simplifies architecture by avoiding hyper-networks and operator prediction modules.
Empirical results in Table 1 demonstrate superior performance of our reconstruction-based approach across multiple benchmarks, with strong generalization capabilities that transfer effectively to unseen transformations. Our contributions are: